# GRADA: Graph-based Reranking against Adversarial Documents Attack

### v2

@@ Abstract
Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3thanks: Corresponding Author., Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu1,511footnotemark: 1 1University of Melbourne, 2University of Edinburgh, 3University College London 4Mohamed bin Zayed University of Artificial Intelligence, 5Macquarie University jingjzheng@student.unimelb.edu.au, zodiachy@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@mbzuai.ac.ae, qiongkai.xu@mq.edu.au
@@ 2 Related Work
When RAG systems came out, Corpus poisoning attacks Zhong et al. (2023) and third-party API attacks Zhao et al. (2024) show a new potential attack surface on LLMs. Later, prompt injection attacks were introduced to bypass the retriever and affect the generator successfully Greshake et al. (2023); Pasquini et al. (2024). However, compared to the prior work, these methods are unstable in the retrieved adversarial passages. While these attacks are based on modern LLM-based retrieval, adversarial manipulation of information-retrieval systems has a much longer history that is instructive for our setting.
@@ 3.2 Reranking
The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval (usually $M\geq n$). In experiments, we retrieve additional documents to maintain consistency in the number of documents in the non-defended scenario.
To prevent adversarial documents from dominating the retrieved set, we ensure that poisoned documents do not constitute the majority by retrieving at least twice the number of documents (i.e., $M\geq 2n$). For instance, if all $n$ original documents are poisoned (e.g., $n=5$), incorporating at least $n$ additional benign documents guarantees that the majority of the final selection is non-poisoned. This approach maintains a substantial presence of benign content in the retrieved set, thereby improving the system’s resilience to adversarial manipulation.
After the algorithm reaches a stationary score distribution, the top $n$ documents are retained, while the remaining documents are discarded. Then, these top $n$ documents are provided as the context of the model.
@@ 4 Experiments
This section begins by detailing the experimental setup, followed by a comparison of our approach with multiple baseline methods. Finally, we analyze the effectiveness of our approach across different settings.
@@ 4.1 Experimental Setup
We conduct experiments on three widely used English datasets: Natural Question Kwiatkowski et al. (2019), MS-MARCO Nguyen et al. (2016) and HotpotQA Yang et al. (2018). The victim models chosen for this study are GPT-3.5-Turbo (version 0125) Brown et al. (2020), GPT-4o (version 2024-08-06) OpenAI et al. (2024), Qwen2.5 Qwen et al. (2025) and LLaMA-3 Grattafiori et al. (2024). The prompts used to generate answers are detailed in Appendix˜A. Contriever Izacard et al. (2021), is a dense retriever model used to find relevant documents by calculating similarity scores between the query and the knowledge base. It was selected for this study due to its efficiency and ability to handle large datasets. In this work, we investigate four distinct attack strategies on RAG. Two of them are Black-box attacks that have no knowledge about the retriever: PoisonedRAG Zou et al. (2024) and PIA Greshake et al. (2023); Pasquini et al. (2024); Perez and Ribeiro (2022). The remaining two are white-box attacks, in which the attacker has access to the victim’s retriever: PoisonedRAG(Hotflip) Zou et al. (2024) and Phantom Chaudhari et al. (2024)
Under default settings without any defense, as in Zou et al. (2024), we retrieve the five most similar documents from the knowledge database to serve as the context for each question. We select 100 close-ended questions from each dataset, yielding 300 questions in total per attack-defense run. Additionally, this process is repeated using 3 random seeds, meaning each attack-defense pair is evaluated on 900 questions in total.
In contrast to Zou et al. (2024), where five poisoned texts are generated and injected into the knowledge base. To provide a more realistic assessment of the attack’s effectiveness, we modify the experiment to inject only a single poisoned document into the database. The original setup, which retrieved only poisoned documents, resulted in a 100% Attack Success Rate (ASR), making it impractical to evaluate the true impact of the attack. As shown in Figure˜3, a similarity cluster of poisoned documents appears in the top-left corner. By applying a clustering algorithm, we can identify and merge redundant information, effectively removing repetitive poisoned entries. This adjustment ensures that only one poisoned document is retrieved, allowing for a more meaningful evaluation of the attack’s success.
@@
We explore three similarity score combinations for GRADA: Embedding-based Distance, BM25, and Hybrid Relevance Similarity with BM25 as the similarity function. [^2] We examine other similarity functions in Section C.2 Here, we utilize Contriever to encode both documents and queries, while for BM25, we adopt the implementation provided by  Lù (2024). We compare GRADA against two reranking models and one defense method: HLATR Zhang et al. (2022), which achieved first place in the MS-MARCO Passage Ranking Leaderboard, BGE-reranker Xiao et al. (2023), which achieves a high precision score in ranking tasks, and Keyword Aggregation Xiang et al. (2024), the only existing defense specifically designed for RAG-based adversarial attacks, as a baseline.
We evaluate the effectiveness of these defense methods by integrating them into our two-stage retrieval system described in Section˜3. We initially retrieve $M=10$ documents, which are then reranked using the aforementioned methods (except for Keyword Aggregation). The top five ranked documents are subsequently provided as the context for the model to answer the query. This ensures that, regardless of the defense configuration, the model always receives a fixed number of five context documents to respond to the question. For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s selection.
@@
Table 2: ASR and EM (%) for various defense methods on White-box attacks.
@@ 4.2 Results and Discussions
As shown in Table˜1, including a single poisoned document in the retrieval process results in a high ASR score. For instance, PoisonedRAG achieves an ASR of 50% across three datasets on both GPT-3.5-Turbo and Llama3.1-8b-Instruct. PIA achieves at least 69% ASR on Llama3.1-8b-Instruct and up to 100% ASR in GPT-3.5-Turbo. These findings emphasize that even minimal adversarial input can achieve very high ASR and degrade the model’s accuracy.
@@
The impact of GRADA on mitigating adversarial attacks is demonstrated in Tables˜1 and 2. As shown in Table˜1, on the NQ and MS-MARCO datasets using GPT-3.5-Turbo, the ASR for PIA decreases from 98.0% and 88.0% to 2.0% and 3.0% by using D2DSIM-EBD. With D2DSIM-EBD, GRADA is also effective against PoisonedRAG, effectively reducing the ASRs from 55.7% and 46.5% to 26.1% and 29.0%. However, the reduction of ASR against PoisonedRAG is more modest than against the other attacks. On this attack, D2DSIM-BM25 and HRSIM led to significant improvements compared to D2DSIM-EBD, where D2DSIM-BM25 achieved an extra 13% decrease in ASR to 13.5% and 16.5%. Beyond that, HRSIM which introduces penalties for excessive similarity to the query, finalizes the ASR to 3% and 8.5%.
@@
Keyword Aggregation is able to reduce ASR significantly, especially for attacks like PIA and Phantom. Keyword Aggregation works by extracting keywords from the answers of each passage to generate the final response, effectively neutralizing attack payloads designed to manipulate or deny answers, such as producing advertisements. However, while it reduces ASR effectively, its EM scores are lower than those of GRADA. For example, on Llama3.1-8b-Instruct in Table˜1, GRADA’s EM scores dominate Keyword Aggregation with at most 21% difference, as some critical information may be lost during keyword extraction. This shows the ability of GRADA to perform well on normal answers even after mitigating adversarial contents.
Similar results to those presented in Table˜1 can also be observed in Table˜2. Notably, GRADA combined with HRSIM consistently outperforms all other approaches, demonstrating that HRSIM is a strong similarity scoring function compared to the alternatives used in GRADA.
@@
GRADA with D2DSIM-EBD effectively preserves model performance on benign inputs across all datasets, with EM score deviations remaining within 4%. Notably, the use of D2DSIM-BM25 leads to a 6% improvement in EM scores on NQ, matching the performance of BGE-reranker, which achieves the highest EM overall. However, on HotpotQA, HRSIM resulted in a 14% reduction in EM scores when handling benign inputs. While this trade-off is significant, it corresponds to HRSIM’s remarkable ASR reduction. Striking a balance between retrieval quality and defense robustness remains a crucial challenge for future research.
Keyword Aggregation has a much lower performance also in EM scores on benign input compared to GRADA. For example, in MS-MARCO, it results in 40% compared to 57% on Llama3.1-8b-Instruct and 59% compared to 75.3% on GPT-3.5-Turbo. Indeed showing the cost of discarding valuable information when facing benign documents.
Using GRADA, we demonstrate that it is possible to defend against the chosen attacks without compromising the model’s overall performance on EM. While reranking methods such as HLATR and BGE-reranker show promise in reducing noise, their limited effectiveness in countering adversarial attack noise highlights a critical gap in existing defenses. Similarly, Keyword Aggregation presents a valuable strategy for mitigating attack payloads but comes with trade-offs in EM scores.
@@
For an attack to be effective, adversaries must ensure that the retriever selects the poisoned documents. To accomplish this, they typically craft these documents to closely resemble the query, exploiting the fact that most retrieval models prioritize query-document similarity. However, these adversarial documents often exhibit only weak similarity to the rest of the corpus, a property that makes them less likely to be flagged by defense mechanisms based on inter-document similarity comparisons.
GRADA leverages this insight by constructing a document similarity graph in which each document effectively "votes" for other documents with which it shares strong semantic similarity. Benign documents, which naturally cluster around shared content, tend to form densely connected subgraphs with high mutual similarity (e.g., averaging $0.82$), thereby reinforcing each other. In contrast, poisoned documents—engineered to deceive—are typically more isolated, receiving fewer "votes" due to their low average similarity to genuine documents (e.g., $0.35$). As a result, GRADA amplifies the collective influence of benign content while attenuating the impact of sparsely connected adversarial documents. A running example is provided in Figure˜13.
Figure 4: Impact of the $\alpha$ value as it increases with three metrics (ASR, number of poisoned documents, and EM) on NQ dataset with GPT-3.5-Turbo.
Impact of $\alpha$ and $M$.
@@
Figure 5: Impact of the $M$ value as it changes with three metrics (ASR, number of poisoned documents, and EM) on MSMARCO dataset with GPT-3.5-Turbo.
@@ 5 Conclusion
The study examines the robustness challenges faced by RAG systems. We identify a critical vulnerability in current adversarial attacks, which focus on increasing semantic similarity to the query without accounting for the relationships between the retrieved documents. Our proposed graph-based filtering framework, GRADA, enhances the robustness of RAG systems by leveraging document similarities and effectively mitigating adversarial impacts through information flow. Experimental results on datasets such as MS-MARCO and NQ, demonstrate at least 30% reductions in ASR across various adversarial strategies. Overall, this work presents a promising direction for developing more secure and reliable RAG systems.

### v3

@@ Abstract
Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3thanks: Corresponding Authors. Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu1,511footnotemark: 1 1University of Melbourne, 2University of Edinburgh, 3University College London 4Mohamed bin Zayed University of Artificial Intelligence, 5Macquarie University jingjzheng@student.unimelb.edu.au, z.xuanli.he@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@mbzuai.ac.ae, qiongkai.xu@mq.edu.au
@@ 2 Related Work
When RAG systems came out, Corpus poisoning attacks Zhong et al. (2023) and third-party API attacks Zhao et al. (2024) show a new potential attack surface on LLMs. Later, prompt injection attacks were introduced to bypass the retriever and affect the generator successfully Greshake et al. (2023); Pasquini et al. (2024). However, compared to the prior work, these methods are unstable in the retrieved adversarial passages. While these attacks are based on modern LLM‑based retrieval, adversarial manipulation of information‑retrieval systems has a much longer history that is instructive for our setting.
@@ 3.2 Reranking
The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval. To prevent adversarial documents from dominating the retrieved set, we ensure that poisoned documents do not constitute the majority by retrieving at least twice the number of documents (i.e., $M\geq 2n$). For instance, if all $n$ original documents are poisoned (e.g., $n=5$), incorporating at least $n$ additional benign documents guarantees that the majority of the final selection is non-poisoned. This approach maintains a substantial presence of benign content in the retrieved set, thereby improving the system’s resilience to adversarial manipulation.
After the algorithm reaches a stationary score distribution, the top $n$ documents are retained, while the remaining documents are discarded. Then, these top $n$ documents are provided as the context to the generative model.
@@ 4 Experiments
This section begins with the experimental setup, followed by a comparison of our approach with multiple baseline methods. Finally, we compare and analyze our approach across different settings.
@@ 4.1 Experimental Setup
We conduct experiments on three widely used English datasets: Natural Question Kwiatkowski et al. (2019), MS-MARCO Nguyen et al. (2016) and HotpotQA Yang et al. (2018). The victim models chosen for this study are GPT-3.5-Turbo (version 0125) Brown et al. (2020), GPT-4o (version 2024-08-06) OpenAI et al. (2024), Qwen2.5 Qwen et al. (2025) and LLaMA-3 Grattafiori et al. (2024). The prompts used to generate answers are detailed in Appendix˜A. Contriever Izacard et al. (2021) is a dense retriever model used to find relevant documents by calculating similarity scores between the query and documents in the knowledge base. It was selected due to its efficiency and ability to handle large datasets. In this work, we investigate four distinct attack strategies on RAG. Two of them are Black-box attacks that have no knowledge about the retriever: PoisonedRAG Zou et al. (2024) and PIA Greshake et al. (2023); Pasquini et al. (2024); Perez and Ribeiro (2022). The remaining two are white-box attacks, in which attackers have the access to the victim’s retriever, i.e., PoisonedRAG (Hotflip) Zou et al. (2024) and Phantom Chaudhari et al. (2024)
Under default settings without defense, as in Zou et al. (2024), we retrieve the five most similar documents from the knowledge database to serve as the context for each question. We select 100 close-ended questions from each dataset, yielding 300 questions in total per attack-defense experiment. Additionally, this process is repeated using 3 random seeds, meaning each attack-defense pair is evaluated on 900 questions in total.
In contrast, Zou et al. (2024) generated five poisoned texts and injected them into the retriever knowledge base. To provide a more realistic assessment of the attack’s effectiveness, we modify the experiment to inject only a single poisoned document into the database. The original setup, which retrieved only poisoned documents, resulted in a 100% Attack Success Rate (ASR), making it impractical to evaluate the true impact of the attack. As shown in Figure˜3, a similarity matrix appears to cluster poisoned documents in the top-left corner. By applying a clustering algorithm, we can identify and merge redundant information, effectively removing repetitive poisoned entries. This adjustment ensures that only one poisoned document is retrieved, allowing for a more meaningful evaluation of the attack performance.
@@
We explore three similarity score combinations for GRADA: Embedding-based Distance, BM25, and Hybrid Relevance Similarity with BM25 as the similarity function. [^2] We examine other similarity functions in Section C.2 Here, we utilize Contriever to encode both documents and queries, while for BM25, we adopt the implementation provided by  Lù (2024). We compare GRADA against two reranking models and one defense method: HLATR Zhang et al. (2022), which achieved first place in the MS-MARCO Passage Ranking Leaderboard, BGE-reranker Xiao et al. (2023), which achieves a high precision score in ranking tasks, and Keyword Aggregation Xiang et al. (2024), the only existing defense specifically designed for RAG-based adversarial attacks.
We evaluate the effectiveness of these defense methods by integrating them into our two-stage retrieval system described in Section˜3. We initially retrieve $M=10$ documents, which are then reranked using the aforementioned methods (except for Keyword Aggregation). The top five ranked documents are subsequently provided as the context for the model to answer the query. This ensures that, regardless of the defense configuration, the model always receives a fixed number of five context documents to respond to the question. For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s keyword selections.
@@
Table 2: Table 2: ASR and EM (%) for various defense methods on the white-box attack methods on GPT-3.5-Turbo and Llama3.1-8b-Instruct.
@@ 4.2 Results and Discussions
As shown in Table˜1, including a single poisoned document in the retrieval process results in a high ASR. For instance, PoisonedRAG achieves an ASR of around 50% across three datasets on both GPT-3.5-Turbo and Llama3.1-8b-Instruct. PIA achieves at least 69% ASR on Llama3.1-8b-Instruct and up to 100% ASR in GPT-3.5-Turbo. These findings emphasize that even minimal adversarial input can achieve very high ASR and degrade the model’s accuracy.
@@
The impact of GRADA on mitigating adversarial attacks is demonstrated in Tables˜1 and 2. As shown in Table˜1, on the NQ and MS-MARCO datasets using GPT-3.5-Turbo, the ASR for PIA decreases from 98.0% and 88.0% to 2.0% and 3.0% by using D2DSIM-EBD. With D2DSIM-EBD, GRADA is also effective against PoisonedRAG, effectively reducing ASRs from 55.7% and 46.5% to 26.1% and 29.0%. However, the reduction of ASR against PoisonedRAG is more modest than against the other attacks. In this attack, D2DSIM-BM25 and HRSIM led to significant improvements compared to D2DSIM-EBD, where D2DSIM-BM25 achieved an extra 13% decrease in ASR to 13.5% and 16.5%. Beyond that, HRSIM which introduces penalties for excessive similarity to the query, finalizes the ASR to 3% and 8.5%.
@@
Keyword Aggregation is able to reduce ASR significantly, especially for attacks like PIA and Phantom. Keyword Aggregation works by extracting keywords from the answers of each passage to generate the final response, effectively neutralizing attack payloads designed to manipulate or deny answers, such as producing advertisements. While Keyword Aggregation reduces ASR effectively, its EM scores are usually lower than those of GRADA. For example, on Llama3.1-8b-Instruct in Table˜1, GRADA’s EM scores dominate Keyword Aggregation with at most 21% difference, as some critical information may be lost during keyword extraction. This shows the ability of GRADA to perform well on normal answers even after mitigating adversarial contents.
Similar results to those presented in Table˜1 can be observed in Table˜2 as well. Notably, GRADA combined with HRSIM consistently outperforms all other approaches, demonstrating that HRSIM is a strong similarity scoring function compared to the alternatives used in GRADA.
@@
GRADA with D2DSIM-EBD effectively preserves model performance on benign inputs across all datasets, with EM score deviations remaining within 3%. Notably, the use of D2DSIM-BM25 leads to a 6% improvement in EM scores on NQ, matching the performance of BGE-reranker, which achieves the highest EM overall. However, on HotpotQA, HRSIM resulted in a 14% reduction in EM scores when handling benign inputs. While this trade-off is significant, it corresponds to HRSIM’s remarkable ASR reduction. Striking a balance between retrieval quality and defense robustness remains a crucial challenge for future research.
Keyword Aggregation has a much lower performance also in EM scores on benign input compared to GRADA. For example, in MS-MARCO, it results in 36% compared to 57% on Llama3.1-8b-Instruct and 59% compared to 75.3% on GPT-3.5-Turbo. Indeed showing the cost of discarding valuable information when facing benign documents.
Using GRADA, we demonstrate that it is possible to defend against the chosen attacks without compromising the model’s overall performance. While reranking methods such as HLATR and BGE-reranker show promise in reducing noise, their limited effectiveness in countering adversarial attack noise highlights a critical gap in existing defenses. Similarly, Keyword Aggregation presents a valuable strategy for mitigating attack payloads but comes with significant trade-offs in EM scores.
@@
For effective attacks, adversaries should steer the retriever to select the poisoned documents. To accomplish this, they typically craft these documents to closely resemble the query, exploiting the fact that most retrieval models prioritize query-document similarity. However, these adversarial documents often exhibit only weak similarity to the rest of the corpus, a property that makes them less likely to be flagged by defense mechanisms based on inter-document similarity comparisons.
GRADA leverages this insight by constructing a document similarity graph in which each document effectively “votes” for other documents with which it shares strong semantic similarity. Benign documents, which naturally cluster around shared content, tend to form densely connected subgraphs with high mutual similarity (e.g., averaging $0.82$), thereby reinforcing each other. In contrast, poisoned documents—engineered to deceive—are typically more isolated, receiving fewer “votes” due to their low average similarity to genuine documents (e.g., $0.35$). As a result, GRADA amplifies the collective influence of benign content while attenuating the impact of sparsely connected adversarial documents. A running example is provided in Figure˜13 in Appendix.
Figure 4: Comparison of $\alpha$ on three metrics (ASR, number of poisoned documents, and EM), based on NQ dataset with GPT-3.5-Turbo.
Impact of hyper-parameters $\alpha$ and $M$.
@@
Figure 5: Comparison of $M$ value on three metrics (ASR, number of poisoned documents, and EM), based on MSMARCO dataset with GPT-3.5-Turbo.
@@ 5 Conclusion
Our research examines the robustness challenges faced by RAG systems. We identify a critical vulnerability in current adversarial attacks, which focus on increasing semantic similarity to the query without accounting for the relationships between the retrieved documents. Our proposed graph-based filtering framework, GRADA, enhances the robustness of RAG systems by leveraging document similarities and effectively mitigating adversarial impacts through information flow. Experimental results on datasets such as MS-MARCO and NQ, demonstrate at least 30% reductions in ASR across various adversarial strategies. Overall, this work presents a promising direction for developing more secure and reliable RAG systems.
@@
Acknowledgement
Aryo Pradipta Gema was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. Giwon Hong was supported by the ILCC PhD program (School of Informatics Funding Package) at the University of Edinburgh, School of Informatics. Xuanli He was funded by an industry grant from Cisco. Pasquale Minervini was partially funded by ELIAI, an industry grant from Cisco, and a donation from Accenture LLP. Qiongkai Xu acknowledges support from 2024 FSE Strategic Startup Grant. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.